# Prioritize Alignment in Dataset Distillation

### v1

@@ 3 Method
To alleviate the information misalignment issue, based on trajectory matching (TM) [1, 11], we propose Prioritizing Alignment in Dataset Distillation (PAD). PAD can also be applied to methods based on matching gradients [54] and distributions [53], which are introduced in Appendix A.1.
@@ 4.2 Main Results
We conduct comprehensive experiments to compare the performance of our method with previous works. As the results presented in Table 1, PAD outperforms previous matching-based methods on three datasets except for the case when IPC=1. When compared with kernel-based methods which use a larger network to perform the distillation, our technique exhibits superior performance in most cases, particularly when the compression ratio exceeds $1\%$. As can be observed, PAD performs relatively better when IPC is high, suggesting our filtering out misaligned information strategy becomes increasingly effective as IPC increases.
@@ 5.5 Generalization to Other Methods
In Section 2, we show that the two filtering modules of PAD can be generalized to existing matching-based DD algorithms. Here, we combine PAD with a more recent DD method that achieves great success on high-resolution datasets, $SRe^{2}L$ [48], to show that PAD has good generalizability. As shown in Table 5(5c), by filtering out misaligned information extracted in the squeeze stage, the performance of $SRe^{2}L$ improves on both small and large IPC settings. This further supports our claim that PAD is beneficial to matching-based methods for performance improvement.
@@ 7 Conclusion
Our alignment strategy could also be applied to methods based on matching gradients and distributions (see Appendix A.1). However, due to the limitation of computing resources, for methods based on matching distributions and gradients, we have only validated our method’s effectiveness on DM [53] and DC [54], and $SRe^{2}L$ [48] (see Table 5(5c), Table 6 and Table 7).
Appendix A Appendix
@@ A.2 Data Scheduler
A.3 Experiment Settings
We use DATM [11] as the backbone TM algorithm and our proposed PAD is built upon. Thus, our configurations for distillation, evaluation, and network are consistent with DATM.

### v3

@@ 3 Method
To alleviate the information misalignment issue, based on trajectory matching (TM) [1, 11], we propose Prioritizing Alignment in Dataset Distillation (PAD). PAD can also be applied to methods based on matching gradients [55] and distributions [54], which are introduced in Appendix LABEL:sec:appendix_dm.
@@ 4.2 Main Results
We conduct comprehensive experiments to compare the performance of our method with previous works. As the results presented in Table 1, PAD outperforms previous matching-based methods on three datasets except for the case when IPC=1. When compared with kernel-based methods, which use a larger network to perform the distillation, our technique exhibits superior performance in most cases, particularly when the compression ratio exceeds $1\%$. As can be observed, PAD performs relatively better when IPC is high, enabling the setting of IPC500 on CIFAR-10 to also achieve lossless performance. This suggests that our filtering out misaligned information strategy becomes increasingly effective as IPC increases. More comparisons can be found in Appendix A.3
@@ 5.5 Generalization to Other Methods
In Section 2, we show that the two filtering modules of PAD can be applied on seminal matching-based DD baselines (DC, DM, MTT) and improve their performances remarkably. To catch up with the latest DD progress, we combine PAD with a more recent DD method that achieves great success on high-resolution datasets, $SRe^{2}L$ [49], to show that PAD can also generalize well on other methods. As shown in Table 5c, by filtering out misaligned information extracted in $SRe^{2}L$’s squeeze and recover stages, the performance of $SRe^{2}L$ improves on both small and large IPC settings. Particularly, PAD’s information extraction filter brings more pronounced improvements to smaller IPCs This further validates PAD’s efficacy in aligning information of dataset distillation and demonstrates that PAD also has decent generalizability on other methods that involve an information-extraction-like or information-embedding-like component.
@@ 7 Conclusion
Our alignment strategy could also be applied to methods based on matching gradients and distributions. However, due to the limitation of computing resources, for methods based on matching distributions and gradients, we have only validated our method’s effectiveness on DM [54] and DC [55], and $SRe^{2}L$ [49] (see Table 5(5c), Table 6 and Table 7).
Appendix
Appendix A Additional Experimental Results and Findings
@@ A.2 Data Scheduler
A.3 Comparison with Knowledge Distillation Methods
Recently, new advancements, represented by RDED [43], in DD have applied knowledge distillation to evaluate distilled datasets and achieved remarkable performance improvements. Specifically, a randomly initialized student model is trained on synthetic data alongside a teacher model. The objective is to minimize the Kullback-Leibler (KL) divergence between the student model’s output and the teacher model’s output on the same batch of synthetic data.
To compare with such a method, we adopt its knowledge distillation strategy during evaluation to ensure a fair comparison. Other experimental settings of PAD are the same. As shown in Table 9, PAD demonstrates superior performances in most of the settings, especially when the compressing ratio is higher (larger IPCs). On small IPCs, the knowledge distillation leads to a drop in PAD’s pure DD performance (Table 1). This is because knowledge from a well-trained teacher exceeds the capacity of small IPCs.
Dataset CIFAR10 CIFAR-100 TinyImageNet
IPC 1 50 500 1 10 50 1 10 50
RDED 23.5±0.3 50.2±0.3 87.2±0.5 19.6±0.3 48.1±0.3 57.0±0.1 12.0±0.1 39.6±0.1 47.6±0.2
PAD 26.8±0.6 62.3±0.5 89.6±0.3 16.4±0.4 50.2±0.3 59.3±.2 10.1±0.3 37.2±0.5 48.5±0.3
Table 9: Comparison between PAD and RDED on CIFAR-10, CIFAR-100, and TinyImageNet. PAD achieves superior performances in 6 out of 9 settings. On larger IPCs, PAD’s advantage is more pronounced.
Appendix B Experimental Settings
We use DATM [11] as the backbone TM algorithm, and our proposed PAD is built upon. Thus, our configurations for distillation, evaluation, and network are consistent with DATM.